OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20 I am completely lost on how to do this!!! I will fan and medal please someone help!

OpenStudy (anonymous):

@ganeshie8 @zepdrix @mathmate

OpenStudy (mathmate):

Complete the square for x and y, then you'll get the equation of the circle with centre (h,k) and radius r as in (x-h)^2+(y-k)^2 = r^2

OpenStudy (mathmate):

@cheetah21 Can you do that or do you need further help?

OpenStudy (anonymous):

I do need further help..sorry im just really bad with math :/

OpenStudy (mathmate):

Start with the given equation, we need to transform it to the standard form of the circle, then we can identify (by comparison) h,k and r.

OpenStudy (anonymous):

Okay how do we do that ?

OpenStudy (mathmate):

The given equation is: x2 + 2x + y2 + 4y = 20 After adding a 1 and a 4 on both sides, we put the equation as (x^2+2x+1) + (y^2+2(2y)+4) = 20 +1 +4 It's still the same equation, but what are in the parentheses are perfect squares. Can you factor the two expressions in parentheses?

OpenStudy (anonymous):

x2 + 2x + y2 + 4y = 20 (x2 + 2x) + (y2 + 4y) = 20

OpenStudy (mathmate):

I already put them in perfect squares, you don't need to simplify them. In fact: (x^2+2x+1) + (y^2+2(2y)+4) = 20 +1 +4 is the same as: (x+1)^2 + (y+2)^2 = 25 You can check to make sure by expanding the squares.

OpenStudy (anonymous):

Oh okay im sorry , i was confused i guess :/ so now do we need to solve in the parentheses or what ?

OpenStudy (mathmate):

Next step is easy: you compare (x+1)^2 + (y+2)^2 = 5^2 with the standard form of the circle with centre (h,k) and radius r (x-h)^2+(y-k)^2 = r^2 and decide what the values of h, k and r are.

OpenStudy (anonymous):

(X-h)^2 + (y - K)^2 = r^2 (x-h)^2+(y-k)^2 = r^2 is that what it should look like right now ?? and how do we figure out values ?

OpenStudy (mathmate):

Compare term by term, for example, (x+1)^2 = (x-h)^2, what would be the value of h to make the right hand side (x+1)^2?

OpenStudy (anonymous):

Would that be 1 also?

OpenStudy (mathmate):

If h=1, then we have (x+1)^2 on the left hand side, and (x-1)^2 , so the two sides don't equal. You need to find another value of h.

OpenStudy (anonymous):

so then that would be 2 ?

OpenStudy (mathmate):

If h=2, then we have (x+1)^2 on the left hand side, and (x-2)^2 , so the two sides don't equal. Remember we are looking for h in: (x+1)^2 = (x-h)^2

OpenStudy (anonymous):

oh 4 ?

OpenStudy (mathmate):

(x+1)^2 = (x-4)^2 (are they equal?)

OpenStudy (anonymous):

(5x)=(25x)

OpenStudy (anonymous):

no wait typo

OpenStudy (mathmate):

There is a positive sign on the left, and a negative sign on the right, so h must be negative to make the two sides match, remembering -(-1)=+1. so putting h=-1 will make (x+1)^2 = (x-(-1))^2=(x+1)^2 or both sides equal.

OpenStudy (mathmate):

Can you try finding k from: (y+2)^2 = (y-k)^2 ?

OpenStudy (anonymous):

okay so then -4

OpenStudy (mathmate):

You would take the square root on both sides, so (y+2) = y-k, can you solve for k?

OpenStudy (anonymous):

(3) = yk 3yk Im sorry if thats a dumb answer i just dont know

OpenStudy (mathmate):

From (y+2) = y-k, subtract y from both sides, so y-y+2 = y-y-k or 2=-k, multiply by -1 to get -2=k, so we have h=-1, k=-2, and r=5 From the standard equation of a circle, (x-h)^2+(y-k)^2=r^2 we therefore found centre = (h,k)=-1,-2, and radius = r =5

OpenStudy (anonymous):

wait so thats the answer to the eqaution ? oh okay , well if i put all the stuff we have come up with together can you tell me if its in the correct order?

OpenStudy (mathmate):

put it up, and perhaps someone else will look at it for you. I have to go now, sorry! :(

OpenStudy (anonymous):

Its okay, thank you for all your help!

OpenStudy (mathmate):

no problem! :)